OpenStudy (anonymous):

Given the function f(x) = log3(x + 4), find the value of f−1(3).

OpenStudy (anonymous):

I need to see how to solve this for a review for an exam please

OpenStudy (anonymous):

Given the function f(x) = log3(x + 4), find the value of f−1(3). f−1(3) = 7 f−1(3) = 12 f−1(3) = 23 f−1(3) = 31

OpenStudy (anonymous):

f(x)=log3 with argument (x+4) what is the value of f^-1(16)

OpenStudy (anonymous):

3 not 16